Subject: Diffusion Morning Report 11/12/12
From: Ted Berg <tberg@stanford.edu>
Date: Mon, 12 Nov 2012 14:14:35 -0800
Mon, 12 Nov 2012 14:14:35 -0800
Good Morning All,
*1. New issues**:
* *2. Ongoing issues:*Tylan Poly rewriting recipes,** Tylannitride  
tested again waiting further instructions after alarms,  working on 
Badger interlocks,*,**particles on back of Epi wafers*,*O2 in Silane? on 
Epi*,*:Thermco nitride shutdown can't pump pump here end of week*,*POCl3 
smells*
*3.  Resolved *:*Thermco Poly2engaged clutch,Replaced FGA cylinders, 
Reset Tylan 1 no overtemp, **,Thermco 1 low growth*-ran a test run all 
flows and temps normal,Epi



Good Morning All,
    1. New issues:
    2. Ongoing issues:Tylan Poly rewriting recipes,  Tylannitride  tested again waiting further instructions after alarms,  working on Badger interlocks,,particles on back of Epi wafers,O2 in Silane? on Epi,:Thermco nitride shutdown can't pump pump here end of week,POCl3 smells
    3.  Resolved :Thermco Poly2engaged clutch,Replaced FGA cylinders, Reset Tylan 1 no overtemp, ,Thermco 1 low growth-ran a test run all flows and temps normal,Epi